### v1

# Probing Visual Language Priors in VLMs

Abstract
Despite recent advances in Vision-Language Models (VLMs), many still over-rely on visual language priors existed in their training data rather than true visual reasoning. To examine the situation, we introduce ViLP, a visual question answering (VQA) benchmark that pairs each question with three potential answers and three corresponding images: one image whose answer can be inferred from text alone, and two images that demand visual reasoning. By leveraging image generative models, we ensure significant variation in texture, shape, conceptual combinations, hallucinated elements, and proverb-based contexts, making our benchmark images distinctly out-of-distribution. While humans achieve near-perfect accuracy, modern VLMs falter; for instance, GPT-4 achieves only 66.17% on ViLP. To alleviate this, we propose a self-improving framework in which models generate new VQA pairs and images, then apply pixel-level and semantic corruptions to form “good-bad” image pairs for self-training. Our training objectives compel VLMs to focus more on actual visual inputs and have demonstrated the effectiveness in enhancing the performance of open-source VLMs, including LLaVA-v1.5 and Cambrian.
1 Introduction
Vision-Language Models (VLMs) have advanced text-image interaction, serving as a cornerstone in bridging the gap between visual and textual data (Achiam et al., 2023; Team et al., 2023). However, a persistent challenge for learning-based models, such as ResNets and CLIPs, is their reliance on strong priors overly learned in the training data, sometimes overlooking visual cues for answering questions (Agrawal et al., 2016; Prabhu et al., 2023). For example, when shown an image of a hexagonal moon (Figure 1), a model might incorrectly assert that “the moon is round" due to strong language priors. Similarly, these models may adhere to learned visual priors (Thrush et al., 2022; Sterz et al., 2024), making it difficult for them to comprehend out-of-distribution visual cues, such as a zebra with atypical spot patterns (Figure 1), where humans would easily discern the difference. This raises an important question: do today’s VLMs still over-rely on learned visual language priors?
To investigate this, we probe the visual language priors of VLMs by constructing Question-Image-Answer (QIA) triplets that deliberately deviate from the training data distribution. Unlike existing benchmarks that typically rely on internet-sourced images, which inadvertently favor the visual language priors embedded in training data, we utilize modern image generation models such as DALL·E-3 (Ramesh et al., 2021) and Flux to synthesize out-of-distribution images, supporting uncommon answers. To further challenge the priors of VLMs, we amplify the language priors in questions by introducing distractor facts, making the questions more suggestive.
Our benchmark, ViLP, contains 300 carefully designed questions, each paired with three distinct answers: a Prior Answer and two Test Answers, resulting in a total of 900 QIA triplets. Each question is structured to present a distractor fact followed by a question. The distractor fact directly leads to the Prior Answer. In contrast, the two Test Answers are crafted to challenge the priors by requiring both textual and visual cues for accurate reasoning. While human participants can differentiate between all three answers, current VLMs face considerable difficulty, as evidenced by a significant performance drop in our benchmarks, with GPT-4o (OpenAI, 2024) scoring only 66.17%.
Inspired by ViLP  results, we propose a self-improving approach to encourage VLM reasoning more relying on visual inputs. Our method involves self-generating VQAs using image generation and editing models (Podell et al., 2023; Ren et al., 2024; Brooks et al., 2023a) and introducing controlled corruptions to create “good" and “bad" image pairs for DPO-like training (Rafailov et al., 2024). We theoretically prove that our proposed objective optimizes an upper bound of the RLHF objective (Ouyang et al., 2022), and experiments on open-sourced VLMs, including LLaVA-v1.5 (Liu et al., 2024) and Cambrian (Tong et al., 2024), demonstrate its effectiveness.
Figure 1: Sample data from ViLP. For the same question, ViLP  provides three distinct images and corresponding answers (upper-left corner). All questions follow a consistent structure, combining a distractor fact with a question. The Prior Answer (first column) can be directly inferred from the question, while Test Answers (second &\&& third column) rely on visual cues. All of our answers are designed to be single words, allowing for evaluation without the need for LLMs. We have developed a robust synonym and plural detection pipeline to ensure that open-ended responses do not interfere with the evaluation process. Please refer to Appendix A for more data samples from ViLP.
2 Related Work
VQA Dataset: There have been long-term efforts to establish Visual Question Answering (VQA) datasets from various perspectives, including general visual question answering  (Agrawal et al., 2015; Gurari et al., 2018; Fu et al., 2023; Liu et al., 2023e; Li et al., 2023a; Yu et al., 2023b; Liu et al., 2024), reading text or charts (Singh et al., 2019a; Mathew et al., 2020; 2021; Masry et al., 2022), complicated reasoning (Lu et al., 2022; 2023), probing compositions (Hudson & Manning, 2019; Ma et al., 2022; Thrush et al., 2022; Hsieh et al., 2023; Li et al., 2024a), studying hallucinations (Rohrbach et al., 2018; Li et al., 2023c; Guan et al., 2023), common sense reasoning (Bitton-Guetta et al., 2023b; a), and more (Majumdar et al., 2024; Sterz et al., 2024). This paper proposed a benchmark to evaluate the visual reasoning abilities of VLMs when the questions, answers and images defy common visual language priors. Our approach follows the balanced dataset design in Goyal et al. (2017), where each question is associated with three answers: the first aligns with language priors, while the latter two deviate from them, requiring reliance on visual cues for correct answering. By leveraging state-of-the-art image generation models, our benchmark creates images that defy priors much more effectively than previous works using internet images, such as Goyal et al. (2017); Tong et al. (2024). Moreover, unlike the “trick” category in Sterz et al. (2024), our process first generates question-answer pairs and then synthesizes images that meet the specified conditions, resulting in more challenging visual reasoning tasks. More comparisons with existing datasets are included in Appendix D.2.
Vision Language Models and Language Priors: Multimodal reasoning is essential for machine intelligence, with vision-language models (VLMs) integrating visual perception, text reasoning, instruction following, and generation for complex tasks (Tan & Bansal, 2019; Li et al., 2019; Kim et al., 2021; Wang et al., 2021b; a; Alayrac et al., 2022; Li et al., 2023b; Chen et al., 2022; Jia et al., 2021; Shen et al., 2021; Singh et al., 2021; Liu et al., 2023c; a; Zhao et al., 2023; Chen et al., 2023; Zhu et al., 2024; Li et al., 2024c; Dai et al., 2023; Li et al., 2024c; Yu et al., 2024; Dai et al., 2024; Deitke et al., 2024). Building on the success of language models (Brown et al., 2020; OpenAI, 2023a; Touvron et al., 2023a; b; Chiang et al., 2023) and pre-trained visual encoders (Radford et al., 2021; Desai & Johnson, 2020; Caron et al., 2021; Chen et al., 2024), many recent advancements leverage vision-language paired datasets (Liu et al., 2024; Tong et al., 2024) to fine-tune intermediate connectors between LLMs and visual encoders (Liu et al., 2024). However, these paired datasets are significantly smaller than the large text corpora used for LLM pre-training (OpenAI, 2023b; Soldaini et al., 2024), and many approaches do not update the visual encoder and LLM parameters during fine-tuning, which can lead to persistent language biases. Specifically, the outputs of VLMs become dominated by the language priors learned from the pre-training corpus, reducing the importance of visual inputs (Thrush et al., 2022; Sterz et al., 2024). This issue becomes even more critical when VLMs are presented with deliberately generated images that challenge these biases, as demonstrated in our work. Previous methods and datasets (Goyal et al., 2016; Agrawal et al., 2017; Dancette et al., 2021; Wu et al., 2022; Ramakrishnan et al., 2018; Gouthaman & Mittal, 2020) have attempted to assess these issues by studying visual backbones and curating datasets using simulators (Johnson et al., 2016) or internet images (Zhang et al., 2016). This paper examines these challenges by creating a novel VQA benchmark, featuring carefully designed questions, fact-based distractors, and rare-distribution answers. Furthermore, we utilize advanced image generation techniques to produce realistic images that subvert common language biases as shown in Figure 1.
Self-Rewarding VLM: Self-rewarding LLM (Yuan et al., 2024) has demonstrated the potential of using LLMs to generate and optimize data through directed reference optimization (Rafailov et al., 2024), leading to self-improvements. This line of research has been extended to VLMs, where new answers are generated and rated for DPO training (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024c; Wang et al., 2024b; a). Our findings resonate with this body of work on self-rewarding VLMs; however, we present two key distinctions: (1) our proposed image-DPO method generates multiple images for a single question-image pair, rather than generating multiple answers and associated ratings; (2) unlike previous methods that focused solely on existing images, both our proposed image-DPO and text-DPO pipelines produce a diverse array of new images with pre-trained image generative models, including SDXL (Podell et al., 2023), GroundedSAM (Ren et al., 2024), and InstructPix2Pix (Brooks et al., 2023a). Specifically, as illustrated in Figure 6, our proposed text-DPO is similar to the self-rewarding VLMs mentioned earlier but incorporates the generation of new images using pre-trained image generation models. In contrast, our image-DPO deliberately corrupts images to create multiple degraded versions, which are used as rejected data during DPO training. This is intended to compel the VLM to rely more on the visual input for accurate generation.
3 ViLP  Benchmark
3.1 Design Principles
"What’s the tall animal with a long neck?" Even without any visual cues, humans readily infer that the answer is "giraffe", leveraging language context and commonsense knowledge. But is it necessarily a giraffe? As shown in the bottom-right of Figure 1, there may be an elephant or a camel with a very long neck and much taller than giraffes. This example highlights a potential issue with current large Vision-Language Models (VLMs). As VLMs are built upon Large Language Models (LLMs) trained on web-scale corpora and are fine-tuned with comparatively much smaller image-text datasets, they may develop a strong bias toward certain visual language priors (especially language) presented in training data and overlook real visual input (Thrush et al., 2022; Sterz et al., 2024). Regarding visual language priors, we focus on strong language priors that confuse VLMs to derive direct answers solely from textual cues. Simultaneously, we aim to create images that are distinctly out-of-distribution to challenge potential visual priors which make the model overlook the critical visual cues such as the unusual spots patterns in the Zebra shown in Figure 1.
To assess the visual reasoning abilities of VLMs against visual language priors, we introduce a new benchmark, ViLP, featuring carefully constructed out-of-distribution question-image-answer (QIA) triplets. These triplets are designed based on the below principles ensuring that VLMs rely on both textual and visual cues, rather than learned priors, to arrive at correct answers.
  • Text-Only Inference: Create questions that, when considered alone, lead to a high-confidence answer based solely on textual information.
  • Visual Influence: When an accompanying image is present, the correct answer differs significantly from the text-only response, sometimes contradicting common sense, ensuring the critical role of visual context in ensuring a high-confidence result.
Mathematically, with $Q$ denoting a question, $I$ denoting an image, $A=\{a_{\text{prior}},\cdots,a_{\text{test}},\cdots\}$ being the set of all possible answers, we use $P(a\mid Q)$ and $P(a\mid Q,I)$ to denote the probability of answer $a$ given the question $Q$ alone, and the probability of answer $a$ given both the question $Q$ and the image $I$, respectively. Here $p$ is a prior model, which could either be human’s cognition prior ($P_{\text{human}}$) or a VLM/LLM’s learned visual language prior ($P_{\theta}$). For constructing our benchmark, we used the following objectives:
Criteria One: Without considering the image, the question $Q$ should lead to a high-confidence answer $a_{\text{prior}}$, where $\delta_{1}$ is a high-confidence threshold. $a_{\text{prior}}$ usually satisfies common knowledge, such as “the moon is round" and “Leonardo da Vinci painted the Mona Lisa" as shown in Figure 1.
$P(a_{\text{prior}}\mid Q)\geq\delta_{1}$ (1)
Criteria Two: With the inclusion of the image $I$, the correct answer becomes $a_{\text{test}}$, where $\delta_{2}$ is another high-confidence threshold. Also, $a_{\text{test}}$ is significantly different from $a_{\text{prior}}$ to ensure that the image has a substantial impact on the answer, where $D$ is a divergence measure (e.g., Kullback-Leibler divergence), and $\delta_{3}$ is a threshold indicating significant difference. For instance, the image in the second-to-last row and column of Figure 1 clearly shows the moon as a hexagon.
$P(a_{\text{test}}\mid Q,I)\geq\delta_{2}\quad\text{and}\quad D\big{(}P(a_{\text{prior}}\mid Q),P(a_{\text{test}}\mid Q,I)\big{)}\geq\delta_{3}$ (2)
Criteria Three: We design answer $a_{\text{test}}$ to be a rare choice, one that is unlikely to be inferred from the question $Q$ alone. In contrast, answer $a_{\text{prior}}$ should be clearly incorrect when considering both the question $Q$ and the image $I$, where $\delta_{4}$ represents a low-confidence threshold. Typically, $a_{\text{test}}$ will be out of context. For example, in the image from the 3rd row and 2nd column of Figure 1, Vincent van Gogh serves as $a_{\text{test}}$, inferred from the image, while Leonardo becomes contradictory after incorporating the additional image condition.
$P(a_{\text{test}}\mid Q)\leq\delta_{4}\quad\text{and}\quad P(a_{\text{prior}}\mid Q,I)\leq\delta_{4}$ (3)
In constructing ViLP, we use human cognition prior $P_{\text{human}}$ as guidance for its design. We evaluated the behavior of the learned priors of VLMs/LLMs by benchmarking $P_{\theta}$ against the human prior $P_{\text{human}}$. This benchmarking process implicitly measuring the divergence between $P_{\text{human}}$ and $P_{\theta}$.
3.2 Question-Image-Answer Generation
As outlined in Criteria Three, we aim for the test answer, $a_{\text{test}}$, to appear highly improbable when considered solely in the context of the question $Q$, yet become a clear and confident choice when paired with the image $I$. This objective presents a significant challenge in finding images $I$, as such images may not follow common knowledge and not naturally exist in the real world or the internet. Recent advancements in generative models offer a solution. By utilizing advanced image generation models such as DALL·E-3 (Ramesh et al., 2021) and Flux, which follow text prompts to synthesize images, we can create images that blend unusual elements. This capability allows us to generate images with the necessary visual coherence to satisfy the specific demands of the question and answer, where we create in a way to disobey language priors. To further ensure the generation of question-image-answers (QIAs) that align with the principles outlined in Section 3.1, our benchmark dataset design incorporates substantial human effort, assisted by state-of-the-art language models such as OpenAI-o1 and Claude-3.5-Sonnet, with additional details provided in Appendix D.1.
Briefly, for each question, we designed three different answers where the first answer $a_{\text{prior}}$ could be inferred solely from the question text. The other two answers were specifically designed against language priors, challenging the model to interpret visual cues despite strong textual priors. During evaluation, we focus on the model’s performance on these second and third answers, i.e., $a_{\text{test}}$. After crafting the initial questions and answers, we leverage GPT-4 to generate text prompts that guide the creation of diverse images at scale. This facilitates the production of varied visual data for the next stage, where high-quality QIA triplets are selected through human review and filtering. Both the text and images are refined as necessary when misalignment or ambiguity are identified. We ensure that the final QIA triplets are clear and easily interpretable by humans, as evidenced by the results of our human evaluation results in Table 2. This iterative process involves several cycles of adjustments to achieve the desired quality. Throughout the process, we encountered two main challenges: (1) generating diverse, out-of-distribution, while reasonable question-answer pairs are difficult, as models like OpenAI-o1 and Claude-3.5-Sonnet tend to converge on a limited set of responses, requiring significant human input to introduce variety; and (2) generating images that against with specific priors using state-of-the-art image generation models is also challenging. In some cases, it was necessary to generate hundreds of images to find just one that accurately matched the question and answer, while avoiding ambiguity.
Table 1: Categorical information.
Type Frequency
Texture 16
Shape 20
Conceptual combinations 276
Hallucinated Components 151
Proverbs 17
As a result, we created a total of 300 questions, each paired with three distinct image-answer combinations, resulting in 900 IQA sets. These questions span a broad range of topics, from low-level visual recognition tasks such as texture and shape, to higher-level visual reasoning involving conceptual combinations, hallucinated components, and proverbs. To reinforce the text priors, we also provide a distractor fact before the question, offering additional context for the prior answer $a_{\text{prior}}$ (the one inferred without the image). Some categorical information is shown on the right Table 1. Most of our data are not confined to a single category; instead, they frequently combine multiple categories to challenge even the most advanced VLMs. For instance, in the moon example (6th-row, 2nd-col), we explore shapes and conceptual combinations, such as a round moon paired with a hexagon versus a hexagonal moon. Similarly, the umbrella example (7th-row, 1st-col) incorporates both conceptual combinations and hallucinated components, blending elements like a plunger and a blender. On average, our data span 1.6 categories per question.
3.3 Dataset Evaluation
All of our questions are designed to be answered with a single word, as this design greatly simplifies evaluation protocol compared to assessing full sentences, without requiring LLMs for evaluation. We explicitly instruct the model to provide a single-word answer, and we evaluate the correctness of each response using a binary system. To ensure a fair evaluation, we devote significant effort to building a comprehensive set of synonyms and plural for each answer to detect other valid alternative answers. This ensures that the model is only penalized for actual errors, not for providing synonymous or alternative correct responses.
4 Image DPO
Inspired by our benchmark results, we propose Image DPO, a self-improvement method for enhancing VLMs’ visual reasoning, featuring a new objective and a data generation pipeline using VLMs themselves and pre-trained image models.
4.1 Objective
Figure 2: Illustration of Image DPO. We construct chosen and rejected pairs by corrupting the image with a set of perturbations while keeping the text (questions and answers) unchanged. This setup encourages that the model more focus on the image to differentiate between the chosen and rejected pairs, as both pairs contain identical textual information. The perturbations applied to the images include semantic editing, Gaussian blurring, and pixelization. More details in Appendix C.2.
Existing approaches for VLM self-improvement usually follow the Direct Preference Optimization (DPO) (Rafailov et al., 2024) used in Large Language Models (LLMs), where the model is trained to distinguish between good and bad answers for a fixed image and question. However, this straightforward adaptation is not vision-centered, as the model can still distinguish good and bad answers from the text alone without needing to analyze the image. In contrast, we propose Image DPO, a vision-focused objective that creates good and bad question-image-answer pairs by corrupting the image while keeping the question and answer unchanged (Figure 2).
Mathematically, given the image $I_{w}$, question $Q$ and the corresponding answer $A$, we generate a corrupted image $I_{l}$ by various image editing operations including gaussian blur, pixelation or semantic editing. With the corrupted image $I_{l}$, the triplet of $Q$, $I_{l}$, and $A$ should form a bad question-image-answer pair compared to the original triplet of $Q$, $I_{w}$, and $A$. We train our model to distinguish between good and bad question-image-answer triplets using the objective function:
$L(\theta,\theta_{\text{ref}})=-\mathbb{E}_{Q,I_{w},I_{w},A\sim S}\left[\log\sigma(\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{w})}-\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{l})})\right]$ (4)
Here, $\theta$ represents the VLM model being trained, $\theta_{\text{ref}}$ is the reference VLM model (typically the previous version of $\theta$), $S$ is the training data consisting of good and bad question-image-answer triplets, $\sigma$ is the sigmoid function, and $\alpha$ is a positive constant.
In Appendix B, we demonstrate that this objective optimizes an upper bound of the RLHF objective (Ouyang et al., 2022), sets it apart from the original DPO objective. Intuitively, since the textual inputs and outputs are identical in both good and bad cases, the gradients of this objective push the model to rely more on the vision branch, driving a shift in gradient direction when processing high-quality images $I_{w}$ compared to corrupted images $I_{l}$, as illustrated in Appendix Figure 9. This behavior encourages the model to focus more on image inputs rather than relying solely on text-based reasoning, thereby enhancing its performance on visually-related tasks.
4.2 Data Generation
Training VLMs requires a large amount of question-image-answer (QIA) triplets, which are often scarce and expensive to collect. In response, we present a scalable data generation pipeline that creates new QIAs from existing datasets by leveraging VLMs and pre-trained image generation and editing models, as illustrated in Appendix Figure 10. Our pipeline is akin to the tool usage of VLMs/LLMs as prior works (Gupta & Kembhavi, 2022; Parisi et al., 2022; Yao et al., 2022; Sur’is et al., 2023), where pre-trained image models are invoked as callable functions. Given a seed image, VLMs are tasked with simultaneously selecting appropriate functions (e.g., image generation or editing models) and generating corresponding instructions. These instructions are then used to produce new images, in addition to the seed image, as illustrated in Figure 11. The same VLMs are then employed to generate QA pairs for these newly created images. Following, we apply the mentioned three types of image corruptions to the generated images, constructing good-bad QIA pairs (i.e., $I_{w}$ and $I_{l}$). Specifically, we employ Stable Diffusion XL (Podell et al., 2023; Rombach et al., 2022) for image generation, and use Instruct-Pix2Pix (Brooks et al., 2023a), and Grounded-SAM (Rombach et al., 2022; Ren et al., 2024) for image editing. All of our seed images are sourced from existing datasets, including COCO (Lin et al., 2014), Text2VQA (Singh et al., 2019b), and Visual Genome (Krishna et al., 2017). Comprehensive details are provided in Appendix C.
5 Experiments
Figure 3: Qualitative examples. We show the results from GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B for some challenging cases.
We introduce ViLP, a benchmark designed to evaluate VLMs’ visual reasoning capabilities in the presence of strong visual language priors, consisting of 300 carefully crafted questions. Each question is paired with three distinct images and corresponding answers, forming one QIAprior (QIA: question-image-answer) and two QIAtest, resulting in a total of 900 unique QIAs. The QIAprior aligns with common knowledge and language priors, making it easy to infer even from text alone. The QIAtest, on the other hand, challenges the VLM’s visual reasoning ability against these priors.
ViLP  includes two evaluation settings: ViLPF (with distractor facts provided) and ViLPP (pure questions without distractor facts). For each setting, we report the average accuracy on QIAtest as our focused benchmark Score, and the accuracy on QIAprior as Prior. Results are on the Table 2.
Is the QIA easy for humans? We first evaluate our benchmark using human study. Humans achieved nearly 100% accuracy on QIAprior and over 98% accuracy on QIAtest, demonstrating that our question-image-answer combinations are non-ambiguous for humans to answer. Despite QIAtest being designed as out-of-distribution examples, humans were able to distinguish them.
Humans performed slightly better on ViLPF-QIAprior when distractor facts were provided, as they could easily identify that these facts aligned with the correct answers. Moreover, Score on ViLPF-QIAtest was marginally lower when facts were introduced, as the distractor facts added some noise and caused minor confusion, although the impact of this noise is relatively small. These findings are consistent with the design principles of our benchmark.
Table 2: Benchmarking on ViLP. Please refer to the left text for symbol definitions. †  indicates the model often fails to follow the instructions.
Model ViLPF ViLPP
Score Prior Score Prior
Baseline
Human 98.33 99.67 98.67 96.67
GPT-4o (text only) 0.0 92.33 0.17 71.33
API call only
GPT-4o 66.17 91.00 56.00 87.67
GPT-4V 57.67 88.33 38.33 85.33
GPT-4o-Mini 57.67 89.00 46.67 84.67
Claude-3.5-Sonnet 70.00 84.33 59.33 86.67
Claude-3-Opus 59.17 74.00 43.00 82.67
Claude-3-Sonnet 48.83 83.67 40.33 81.33
Claude-3-Haiku 43.67 82.67 34.83 82.33
Gemini-1.5-Pro 60.50 79.33 48.00 83.00
Gemini-1.5-Flash 54.50 83.33 69.17 79.67
Open weights
Llama-3.2-Vision-11B 67.33 76.67 61.17 79.33
Llama-3.2-Vision-90B 64.00 91.67 63.17 83.33
MolmoE-1B 48.67 57.33 47.83 69.00
Molmo-7B-O 57.83 60.67 47.33 76.33
Molmo-7B-D 54.5 69.00 46.17 72.33
Molmo-72B 60.33 85.00 47.17 82.33
Qwen2-VL-7B 50.50 83.00 48.67 80.33
Qwen2-VL-72B 56.50 92.33 53.83 83.00
InternVL2-8B 47.00 66.67 43.00 75.00
InternVL2-76B 42.67 47.67 50.84 74.33
LLaVA-1.5-7B 29.67 71.33 37.67 65.67
LLaVA-1.5-13B 35.33 81.00 41.50 73.67
Cambrian-1-8B 8.67 43.67 32.50 63.67
LLaVA-OneVision-7B 54.17 82.33 49.67 75.00
LLaVA-OneVision-72B 1.67 3.00 5.22 11.67
Figure 4: Before and after removing distactor facts.
Are our QIAs aligned with the learned priors of VLMs? We evaluated GPT-4o’s performance on our questions without providing any images, i.e., GPT-4o (text only) in Table 2. Additionally, we eliminate image-related phrases from the prompt during this text-only inference, such as removing expressions like “as shown in the image". Remarkably, even without any visual content, the model was able to correctly answer 92.33% of QIAprior when distractor facts were included, demonstrating that our questions are highly suggestive for VLMs, allowing them to answer even without visual input. Once removing the distractor facts, the performance decrease to 71.33%, as these facts implicitly pointed to the correct answers in the fact statements. It is noteworthy that some powerful models like Claude-3-Opus and several open-source VLMs fail to outperform GPT-4o (text-only) performance on QIAprior, even when provided with images. For QIAtest, the success rate of GPT-4o (text-only) dropped dramatically to near zero (0% & 0.17%), indicating that QIAtest are far outside the distribution typically encountered by VLMs or LLMs. These observations further validate the design principles of our benchmark: QIAprior can be easily inferred using text alone, while QIAtest is specifically designed to challenge language-based reasoning, requiring reliance on visual cues to arrive at correct answers.
How do VLMs perform on our benchmark? Although our benchmark questions are distinguishable for humans, they are challenging for VLMs. Even the most advanced VLM models like GPT-4o, have a clear performance gap (more than 25%) compared to humans’ performance on QIAtest (Score in ViLPF), indicating the difficulty of these questions for VLMs. GPT-4o achieves above 91% accuracy on QIAprior (Prior), which is quite close to human performance. While its Score on QIAtest is still low (66.17% versus 98.33%), demonstrating the difficulty of these questions for VLMs. We also provide some qualitative examples in Figure 3, showcasing results from both leading commercial and open-source models, including GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B. They encounter challenges when tackling these cases of our ViLP. Notably, it is encouraging to see that some open-source models achieved over 60% accuracy on ViLPF-QIAtest, with performance nearing that of their commercial counterparts, including Llama-3.2-Vison and Molmo-72B.
Do distractor facts really distract? In ViLPF setting, we add a distractor fact before the question and explicitly instruct the model to answer with one word. Since these facts implicitly suggest incorrect answers for QIAtest, we expected this change to make the questions more suggestive and lower the ViLPFScore, as the distractors would mislead the VLMs.
But is this true in practice? Interestingly, we find that strong models like GPT-4o are not misled by these distractor facts, despite their performance is still significantly lagged behind human accuracy. Not only do these models avoid being misled, but surprisingly, they also leverage the distractor factsto improve their answers, leading to a substantial and consistent performance boost when the facts are included. We illustrate some examples of this behavior in Figure 4. When the distractor fact is provided, GPT-4o performs better on QIAtest compared to when it is not. This result is counterintuitive, as the model performs better when its input prompt includes misleading information. We hypothesize that the distractor fact may help the model infer the focus of the question, thereby narrowing the search space and ultimately leading to higher accuracy.
For models with reasonable instruction-following abilities but limited visual reasoning, like LLaVA-1.5-13B (Liu et al., 2023b), the distractor facts behave as expected, often leading the model to incorrect conclusions based solely on textual information. This leads to a significant performance drop on ViLPF-Score when distractor facts are present, simultaneously boosting performance on ViLPF-Prior, the language-prior answer. For instance, as shown in Figure 4, with including distractor facts, LLaVA-1.5-13B consistently predicts the distractor fact as the answer. However, once the distractors are removed, it can then make the correct prediction.
In contrast, for models with weaker instruction-following abilities, like Cambrian-8B (Tong et al., 2024), the inclusion of distractor facts significantly increases the difficulty of adhering to explicit instructions, such as providing a single-word answer. Specifically, when facts are provided, Cambrian-8B fails to follow the instruction in 62% of the questions, compared to 30% without facts (nearly a 2x increase). Upon manual review of these failures, we find that approximately 59% of the responses that deviate from the instruction are still contextually correct, leading to an adjusted accuracy of 47.92%. A similar issue is observed with LLaVA-OneVision-72B (Li et al., 2024b), which consistently generates sentences with analysis, when explicitly prompted with “please answer this question with one word". This pattern highlights a concerning trend: focusing on improving performance on well-established benchmarks may come at the cost of basic instruction-following abilities, ultimately limiting the practical utility of these models in real-world applications.
Figure 5: Comparison of benchmark scores under different image transformations. Solid line and dotted line refer to ViLPF-Score and ViLPF-Prior, respectively.
How image transformations affect the results? We also investigate how image transformations, including resizing, Gaussian blur, and pixelation, affect ViLPperformance. The results, shown in Figure 5, reveal that the ViLPF-Score rapidly decreases as the severity of the transformations (x-axis) increases, while the ViLPF-Prior score remains around 50%. Interestingly, GPT-4o, when using degraded images, performs worse in ViLPF-Prior than when no images are used, i.e., GPT-4o (text only) in Table 2.
5.1 Image DPO
In this section, we examine our proposed Image-DPO (Section 4.1) on both the ViLP  benchmark and general VQA benchmarks. Specifically, for self-generating question-image-answer (QIA) pairs, we use COCO (Lin et al., 2014), Text2VQA (Singh et al., 2019b), and Visual Genome (Krishna et al., 2017) as our seed datasets. Based on these seed datasets, we instruct the VLM to leverage pretrained image generation and editing models (Liu et al., 2023d; Podell et al., 2023; Brooks et al., 2023b) to generate new images and corresponding QA pairs. After generating the QA pairs, we corrupt the images either semantically or at the pixel level, while keeping the QA pairs unchanged, to create the good and bad QIA pairs used for Image DPO training. Details of our image generation method are provided in the Appendix C and training details are in the Appendix C.2
Table 3: Comparisons of Image-DPO on ViLP.
Model ViLP F ViLP P
Score Prior Score Prior
LLaVA-1.5-7B 29.67 71.33 37.67 65.67
LLaVA-1.5-7B + CSR 27.50 64.33 40.84 63.67
LLaVA-1.5-7B + RLHF-V 29.50 75.00 36.33 65.33
LLaVA-1.5-7B + Text-DPO 31.34 71.67 37.83 65.67
LLaVA-1.5-7B + Image-DPO 34.17 77.67 38.67 67.33
For comparison, we introduce another baseline, Text-DPO, which follows the same question-image-answer generation process as our Image-DPO but applies LLM self-rewarding techniques as outlined in (Yuan et al., 2024). In Text-DPO, good and bad pairs are created by sampling positive and negative answers generated by the VLM, while keeping the image and question unchanged. The difference between Image-DPO and Text-DPO is illustrated in Figure 6. The green box corresponds to Image-DPO, where multiple corrupted images are generated through semantic alterations, Gaussian blurring, and pixelation, while the question and answer remain constant. In contrast, Text-DPO (purple-box) maintains the same image and generates multiple answers, each associated with a rating. It is worth noting that while Text-DPO shares similarities with other VLM self-rewarding techniques (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024c; Wang et al., 2024b; a), a key distinction is our use of pre-trained image generation models to create diverse images.
Table 4: Effectiveness of Image-DPO on General VQA benchmarks. The proposed Image-DPO could consistently improve the performance of various VLMs across general VQA benchmarks.
[…]
Cambrian-8B 8.67 32.50 44.6 47.9 19.4 14.5 4.7 51.4 11.3
Cambrian-8B + Image-DPO 11.50 $\uparrow$ 2.83 35.17 $\uparrow$ 2.67 45.68 $\uparrow$ 1.08 49.45 $\uparrow$ 1.55 20.11 $\uparrow$ 0.71 14.3 $\uparrow$ -0.2 4.8 $\downarrow$ 0.1 52.7 $\uparrow$ 1.3 12.0 $\uparrow$ 0.7
LLaVA-1.5-7B 29.67 37.67 37.7 43.8 12.7 49.1 14.8 31.1 58.6
LLaVA-1.5-7B + Image-DPO 34.17 $\uparrow$ 4.5 38.67 $\uparrow$ 1 40.68 $\uparrow$ 2.98 46.29 $\uparrow$ 2.49 14.95 $\uparrow$ 2.25 45.4 $\uparrow$ -3.7 12.4 $\uparrow$ -2.4 33.8 $\uparrow$ 2.7 59.65 $\uparrow$ 1.05
LLaVA-1.5-13B 35.33 41.50 39.6 44.6 14.8 48.30 14.10 36.10 61.9
LLaVA-1.5-13B + Image-DPO 38.17 $\uparrow$ 2.84 42.5 $\uparrow$ 1 42.68 $\uparrow$ 3.08 47.37 $\uparrow$ 2.77 17.16 $\uparrow$ 2.36 42.6 $\uparrow$ -5.7 11.6 $\uparrow$ -2.5 37.6 $\uparrow$ 1.5 62.32 $\uparrow$ 0.42
Figure 6: Image-DPO v.s. Text-DPO.
For comparison, we also report the results of CSR (Zhou et al., 2024b), a VLM self-improvement method that similarly constructs good and bad answers to form training data for DPO. Specifically, we use the publicly available checkpoint of the CSR model. Additionally, we train a model using RLHF-V (Yu et al., 2023a), a supervised fine-tuning method based on human-annotated data. All models are trained on the LLaVA-7B architecture due to computational constraints. The results, presented in Table 3, show that our Image-DPO method achieved the highest performance in ViLPF-Score, outperforming Text-DPO, CSR (Zhou et al., 2024b), and the RLHF-V model based on human-annotated data (Yu et al., 2023a). On ViLPP-Score, the Image-DPO method achieved the second-best performance.
Besides, we evaluate the proposed Image-DPO algorithm across three VLM models—Cambrian-8B, LLaVA-1.5-7B, and LLaVA-1.5-13B—using several popular VLM benchmarks that focus on different aspects, including compositionality(NaturalBench (Li et al., 2024a)) hallucinations (CHAIR (Rohrbach et al., 2018)), visual conversation (MM-Vet (Yu et al., 2023c)), and visual reasoning, SEED-Bench (Li et al., 2023a)). The results, presented in Table 4, show consistent performance improvements across both datasets and models, further demonstrating the effectiveness of our Image DPO method.
6 Conclusion
In conclusion, we present the ViLPbenchmark to probing the challenge of language bias in Vision-Language Models. By utilizing advanced image generation models and designing questions that demand visual cues for accurate responses, our benchmark consists of images that challenge language priors, revealing the limitations of current VLMs, which tend to rely heavily on text. Our method, which incorporates self-generated VQA pairs and image corruption for training, has shown promising improvements in enhancing visual reliance, as evidenced by performance gains on models like LLaVA-v1.5 and Cambrian.
7 Acknowledgment
This work was supported by the LG AI Research grant. We also extend our gratitude to the OpenAI Researcher Access Program for providing the credits used to access OpenAI’s APIs. Finally, we greatly appreciate Chris Rockwell for his contributions to our human evaluation studies.
Appendix
[…]
Appendix A More data samples of ViLP
Figure 7: Randomly sampled data from ViLP.
Figure 8: Randomly sampled data from ViLP.
Appendix B Image-DPO Mathematical Details
Figure 9: Gradients difference between Image-DPO and Text-DPO. For Text-DPO (Rafailov et al., 2024), the model receives positive gradients (green arrows) for the preferred answer $A_{w}$ and negative gradients $A_{l}$ (red arrows) for the dispreferred answer. In contrast, our proposed Image-DPO approach applies positive gradients when the preferred image $I_{w}$ is input and negative gradients for the dispreferred image $I_{l}$, both based on the same output answer.
In this section, we give the complete proof of Image DPO. Simlarly to DPO, we start from the objectiveness of RLHF and then derivate its variant for image dpo.
B.1 RLHF for VLM
SFT Given question $Q$, answer $A$ and image $I$, we can train a SFT model $\pi^{SFT}$ with supervised learning on high-quality data. As the SFT model is a language generation model, it is still a model modeling the text outputs with question and image. $\pi^{SFT}(A|Q,I)$
Reward Modeling Phase In this stage, we construct a static dataset of comparisons $\mathcal{S}=\{A^{i},Q^{i},I_{w}^{i},I_{l}^{i}\}$, and we present the QIA pairs $(Q,I_{w},A)$, $(Q,I_{l},A)$ to human for preference.
Following the idea of RLHF, the preference are assumed to be obtained from a a latent reward function $r^{*}(Q,I,A)$ which are not tractable, and we we use BT model to represent the preference distribution $p^{*}$ as:
$p^{*}((Q,I_{w},A)\succ(Q,I_{l},A))=\frac{\exp\left(r^{*}\left(Q,I_{w},A\right)\right)}{\exp(r^{*}(Q,I_{w},A)+\exp(r^{*}(Q,I_{l},A))}$ (5)
Now given the human labeled preference, we can try to optimize a reward model $r_{\phi}$ to estimate $r^{*}$ by using maximum likelihood. Framing this as a binary classification, we can have this negative log-likelihood loss:
$\mathcal{L}_{R}(r_{\phi},\mathcal{S})=-\mathbb{E}_{(A,Q,I_{w},I_{l})\sim\mathcal{S}}\left[\log\sigma(r_{\phi}(Q,I_{w},A)-r_{\phi}(Q,I_{l},A))\right]$ (6)
Here $\sigma$ is a logistic function. Basically, this reward function gives score jointly considering image, question and image quality.
RL Fine-Tunning Phrase
During the RL phase, the learned reward function is used to provide feeback to the VLM model. Following DPO paper, the optimization is formulated as :
$\max_{\pi_{\theta}}\mathbb{E}_{(Q,I)\sim\mathcal{S},A\sim\pi_{\theta}(A|Q,I)}\left[r_{\phi}(Q,I,A)\right]-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(A|Q,I)\|\pi_{\text{ref}}(A|Q,I)\right],$ (7)
where $\beta$ is a parameter controlling the deviation from the base reference policy $\pi_{\text{ref}}$, namely the initial SFT model $\pi^{\text{SFT}}$. Due to the discrete nature of language generation, this object is also not differentiable and is typically optimized with reinforcement learning.
B.2 Image DPO and RLHF
According to the DPO paper, a straightforward optimal solution to the KL-constrainted reward function maximization object in Eq. 6 is:
$\pi_{r}(A|Q,I)=\frac{1}{Z(Q,I)}\pi_{\text{ref}}(A|Q,I)\exp(\frac{1}{\beta}r(Q,I,A))$ (8)
where $Z(Q,I)=\sum_{A}\pi_{\text{ref}}(A|Q,I)\exp(\frac{1}{\beta}r(Q,I,A)$ is a partition function. Here $r$ should be any reward function, which makes $Z$ hard to tract. We provide the proof of this step in B.3.
Taking the logarithm of both side, and with some algebra, we get
$r(Q,I,A)=\beta\frac{\pi_{r}(A|Q,I)}{\pi_{\text{ref}}(A|Q,I)}+\beta\log Z(Q,I)$ (9)
This parametrization could be applied to ground-truth reward $r^{*}$ and the corresponding optimal model $\pi^{*}$.
The BT model with the optimal policy is
$p^{*}((Q,I_{1},A)\succ(Q,I_{2},A))=\frac{\exp\left(r^{*}(Q,I_{w},A)\right)}{\exp\left(r^{*}(Q,I_{w},A)\right)+\exp\left(r^{*}(Q,I_{l},A)\right)}$ (10)
We plug Eq. 9 into the BT model, we have:
$p^{*}((Q,I_{1},A)\succ(Q,I_{2},A))$ $=\frac{\exp\left(\beta\log\frac{\pi^{*}(A|Q,I_{w})}{\pi_{\text{ref}}(A|Q,I_{w})}+\beta\log Z(Q,I_{w})\right)}{\left(\beta\log\frac{\pi^{*}(A|Q,I_{w})}{\pi_{\text{ref}}(A|Q,I_{w})}+\beta\log Z(Q,I_{w})\right)+\left(\beta\log\frac{\pi^{*}(A|Q,I_{l})}{\pi_{\text{ref}}(A|Q,I_{l})}+\beta\log Z(Q,I_{l})\right)}$
$=\frac{1}{1+\exp(\beta\log\frac{\pi^{*}(A|I_{l},Q)}{\pi_{\text{ref}}(A|I_{l},Q)}-\beta\log\frac{\pi^{*}(A|I_{w},Q)}{\pi_{\text{ref}}(A|I_{w},Q)}+\beta\log Z(I_{l},Q)-\beta\log Z(I_{w},Q))}$
$=\sigma\left(\exp(\beta\log\frac{\pi^{*}(A|I_{l},Q)}{\pi_{\text{ref}}(A|I_{l},Q)}-\beta\log\frac{\pi^{*}(A|I_{w},Q)}{\pi_{\text{ref}}(A|I_{w},Q)}+\beta\log Z(I_{l},Q)-\beta\log Z(I_{w},Q))\right)$
Now we have the probability of human preference data in terms of the optimal policy rather than the reward model, we can formulate a maximum likelihood objective for a policy $\pi_{\theta}$. Our policy objective is :
$\mathcal{L}(\pi_{\theta};\pi_{\text{ref}})=\mathbb{E}_{(Q,A,I_{w},I_{l})\sim\mathcal{S}}\left[-\log\sigma\left(\beta\log\frac{\pi_{\theta}(A|I_{w},Q)}{\pi_{\text{ref}}(A|I_{w},Q)}-\beta\log\frac{\pi_{\theta}(A|I_{l},Q)}{\pi_{\text{ref}}(A|I_{l},Q)}+\beta\log Z(I_{w},Q)-\beta\log Z(I_{l},Q)\right)\right]$ (11)
As $f(x)=-\log\sigma(x)$ is a convex function ($\sigma$ is the sigmoid function), we can apply Jensen’s inequality $f(\frac{1}{2}x+\frac{1}{2}y)\leq\frac{1}{2}f(x)+\frac{1}{2}f(y)$:
$\mathcal{L}(\pi_{\theta};\pi_{\text{ref}})\leq\mathbb{E}\left[-\frac{1}{2}\log\sigma\left(2\beta\log\frac{\pi_{\theta}(A|I_{w},Q)}{\pi_{\text{ref}}(A|I_{w},Q)}-2\beta\log\frac{\pi_{\theta}(A|I_{l},Q)}{\pi_{\text{ref}}(A|I_{l},Q)}\right)-\frac{1}{2}\log\sigma\left(2\beta\log Z(I_{w},Q)-2\beta\log Z(I_{l},Q)\right)\right]$ (12)
As $\log\sigma\left(Z(I,Q)\right)$ is not a function of $\pi_{\theta}$, the above objective is equivalent to the below Eq.13, where $\alpha=2\beta$. It is the same as our objective listed in Eq.4 of the main paper.
$\mathcal{L}(\pi_{\theta};\pi_{\text{ref}})\leq-\mathbb{E}_{(Q,I_{w},I_{l},A)\sim S}\left[\log\sigma(\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{w})}-\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{l})})\right]$ (13)
In this sense, our optimization objective Eq.4 in main paper are optimizing the upper bound of RLHF, i.e., Eq.7.
B.3 Deriving the Optimum of the KL-Constrained Reward Maximization Objective
In this appendix, we will derive Eq.8. Similarly to Eq.7, we optimize the following objective:
$\max_{\pi}\mathbb{E}_{(Q,I)\sim\mathcal{S},A\sim\pi}\left[r(Q,I,A)\right]-\beta D_{\mathrm{KL}}\left[\pi(A|Q,I)\|\pi_{\mathrm{ref}}(A|Q,I)\right]$ (14)
under any reward function $r(Q,I,A)$, reference model $\pi_{\mathrm{ref}}$, and a general non-parametric policy class. We now have:
$\max_{\pi}\mathbb{E}_{(Q,I)\sim\mathcal{S},A\sim\pi}\left[r(Q,I,A)\right]-\beta D_{\mathrm{KL}}\left[\pi(A|Q,I)\|\pi_{\mathrm{ref}}(A|Q,I)\right]$ (15)
$=\max_{\pi}\mathbb{E}_{(Q,I)\sim\mathcal{S}}\mathbb{E}_{A\sim\pi(A|Q,I)}\left[r(Q,I,A)-\beta\log\frac{\pi(A|Q,I)}{\pi_{\mathrm{ref}}(A|Q,I)}\right]$
$=\min_{\pi}\mathbb{E}_{(Q,I)\sim\mathcal{S}}\mathbb{E}_{A\sim\pi(A|Q,I)}\left[\log\frac{\pi(A|Q,I)}{\pi_{\mathrm{ref}}(A|Q,I)}-\frac{1}{\beta}r(Q,I,A)\right]$
$=\min_{\pi}\mathbb{E}_{(Q,I)\sim\mathcal{S}}\mathbb{E}_{A\sim\pi(A|Q,I)}\left[\log\frac{\pi(A|Q,I)}{\frac{1}{Z(Q,I)}\pi_{\mathrm{ref}}(A|Q,I)\exp\left(\frac{1}{\beta}r(Q,I,A)\right)}-\log Z(Q,I)\right]$
where we have the partition function:
$Z(Q,I)=\sum_{A}\pi_{\mathrm{ref}}(A|Q,I)\exp\left(\frac{1}{\beta}r(Q,I,A)\right)$ (16)
Observe that the partition function depends solely on $(Q,I)$ and the reference policy $\pi_{\mathrm{ref}}$, and is independent of the policy $\pi$. We can now define the Equation 8.
$\pi^{*}(A|Q,I)=\frac{1}{Z(Q,I)}\pi_{\mathrm{ref}}(A|Q,I)\exp\left(\frac{1}{\beta}r(Q,I,A)\right),$ (17)
Appendix C Details in Image-DPO data generation and training
Our image-DPO data generation pipeline consists of two stages. In the first stage, we leverage the VLM we aim to enhance to perform self-guided data generation with the aid of pre-trained image generative models. This stage produces a large number of new question-image-answer (QIA) triplets. In the second stage, we apply three types of image corruptions—Gaussian blurring, pixelization, and semantic editing—to generate good-bad QIA pairs, denoted as $I_{w}$ (good) and $I_{l}$ (bad).
Details of the hyperparameters used in the experiments are provided at the end of this section.
C.1 VLM self-guided data generation
Figure 10: Overview of our data generation pipeline. We begin with an image only, from which instructions are derived using a VLM. These instructions guide the creation of a new image or the modification of the existing one. The generated image is then processed by the VLM to generate QA pairs. Both the QA pair and the image are subsequently input back into the VLM to assess the quality of the answers. No human-written in-context examples are used throughout this process.
As illustrated in Figure 10, our data generation process begins by utilizing VLMs to suggest modifications or draw inspiration for input images without relying on any in-context examples. The used text prompt is shown in Figure 12. Subsequently, pre-trained models such as Stable Diffusion XL Podell et al. (2023), Instruct-Pix2Pix Brooks et al. (2023b), and Grounded-SAM Ren et al. (2024) are employed to either modify existing images or generate entirely new ones.
The altered or newly created images, along with the instructions that guided their generation, are then used by the same VLMs to produce corresponding question-answer pairs (QAs) based on the text prompt shown in Figure 13. An example of this process is provided in Figure 11. Importantly, all instructions, tool selections, and QA generation are autonomously handled by the same VLM we aimed to improve.
In particular, Grounded-SAM requires the VLM to specify the object to be modified before generating images. To facilitate this, we use an additional text prompt (Figure 14) after the VLM generates the initial instructions (the pink region of Figure 11).
To provide a better understanding of our generated QIAs, we randomly sampled and listed some examples of the generated QIA data, as shown in Figures 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, and 29.
Figure 11: Randomly sampled generation examples. Our proposed data generation pipeline utilizes pretrained model to generate diverse new images from existing seed image datasets.
Given this image, please suggest a range of creative edits, tasks, or transformations that could be applied using advanced image processing tools. These tasks may include artistic transformations, vivid color adjustments, object detection and modification, or completely creating a new image inspired by the original. Specify which tool would be best suited for each task, choosing from Stable Diffusion for image generation, InstructPix2Pix for image modification, or GroundingDINO for object modification. Your recommendations should help in understanding the potential of the image and exploring creative possibilities. Expected Response Format: Item Number: 1 Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO] Text Prompt for Processing: [Detailed description of the task or transformation to be performed. For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.] Item Number: 2 Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO] Text Prompt for Processing: [Detailed description of the task or transformation to be performed. For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.] Item Number: 3 Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO] Text Prompt for Processing: [Detailed description of the task or transformation to be performed. For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.]
Figure 12: The prompt for instruction generation. We ask the VLM to generate instructions for using pre-trained image models.
Given this image, could you please generate a series of insightful and diverse question-answer pairs based on the image and its descriptions? We are interested in exploring various facets of the image, including: - Holistic styles and layouts: Questions that analyze the overall design, style, and layout of the image. - Object-specific details: Questions that delve into particular elements or objects within the image, discussing their characteristics or functions. - Background context: Questions that speculate about the background story or the setting of the image. - Overall themes: Questions that interpret the thematic elements and messages portrayed in the image. We encourage creative and thought-provoking questions that extend beyond the basics. Please generate questions that cover a broader range of observations and insights drawn from the image. Each question should be followed by a comprehensive answer, providing depth and context. Expected Multiple Response Format: Item Number: 1 Question: [Propose a unique and insightful question based on the descriptions and the images.] Answer: [Provide a comprehensive answer to the proposed question.] Item Number: 2 Question: [Propose a unique and insightful question based on the descriptions and the images.] Answer: [Provide a comprehensive answer to the proposed question.] Please ensure each question-answer pair is well-defined and informative. Please provide at least 5 question-answer pairs based on the input provided.
Figure 13: The prompt for single-image QAs. We ask the VLM itself to generate single-image QAs based on the generated images by pre-trained models.
Analyze the provided image and its accompanying modification instruction to identify the removed object description, the new object description, and the new image description. Modification Instructions: <Text Prompt for Processing> Expected Multiple Response Format: Item Number: 1 Removed Object Description: [Brief description of the object to be detected and removed] New Object Description: [Description of a new, different object to replace the removed one] New Image Description: [Description of the image after each object’s removal, focusing on changes and remaining elements] Item Number: 2 Removed Object Description: [Brief description of the object to be detected and removed] New Object Description: [Description of a new, different object to replace the removed one] New Image Description: [Description of the image after each object’s removal, focusing on changes and remaining elements]
Figure 14: The prompt for instruction generation of Grounded-SAM. We ask the VLM to generate designated instructions to use Grounded-SAM.
C.2 Image DPO data preparation and training details
This section details the construction of good-bad question-image-answer (QIA) pairs ($I_{w}$, $I_{l}$) based on the QIAs generated by the pipeline described in Appendix C.1. In brief, the data generation pipeline outlined in Appendix C.1 utilizes VLMs in conjunction with pre-trained image models to generate or modify images and create corresponding question-answer pairs. This process results in a collection of QIA triplets, as illustrated in the Figures 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, and 29.
After generating the QIA triplets, we apply three image corruption methods—Gaussian blurring, pixelization, and semantic editing—to create good-bad QIA pairs for ImageDPO training (Section 4.1), while keeping the QA components unchanged.
For Gaussian blur, we use a kernel size of 40 for Cambrian and 80 for LLaVA, as the larger kernel size showed better performance for LLaVA. For pixelization, we apply block sizes of 32 for Cambrian and 64 for LLaVA. For these two types of corruption, we utilize data generated by all three pre-trained models: Stable Diffusion, Instruct-Pix2Pix, and Grounded-SAM.
However, for semantic editing, we rely solely on data from Grounded-SAM, which modifies a single object in the image while leaving the rest unchanged. The object to be altered and the type of modification are determined by the VLMs based on the text prompt (Figure 14). For instance, the VLM decides to add a headset to the chef, as shown in Figure 15. Grounded-SAM then modifies the image accordingly by adding the headset, after which the VLM generates a single QA pair based on the text prompt (Figure 13). Note that, we retain only images where the altered region covers more than 10% of the image size, as this ensures a significant and distinct change.
For model training, we constructed datasets comprising 800k and 400k DPO pairs to fine-tune LLaVA (7B and 13B) and Cambrian-8B, respectively. Batch sizes are set to 112 for LLaVA-7B, 80 for LLaVA-13B, and 8 (with 4 gradient accumulation steps) for Cambrian-8B. We employ LoRA with a rank of 128, an alpha of 256, and a learning rate of 5e-7, training each model for 2 epochs. The GPUs we used are 8-L40S.
Figure 15: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 16: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 17: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 18: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 19: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 20: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 21: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 22: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 23: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 24: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 25: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 26: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 27: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 28: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 29: Randomly sampled single-image QA generated by our data generation pipeline.
Appendix D More details and comparisons of our benchmarks
D.1 Our benchmark data generation
Our proposed dataset introduces QIA triplets meticulously designed to challenge state-of-the-art Vision-Language Models (VLMs) by mitigating language priors. The construction process combines human-guided and automated efforts to ensure quality and alignment across all components.
Question-Answer (QA) Generation: Questions and answers are partially authored by humans, adhering to the design principles outlined in Section 3.1. Additionally, candidate QA pairs are generated using models like OpenAI-O1 and Claude-3.5-Sonnet with carefully crafted prompts. One example prompts shown in Figure 30. These candidates undergo rigorous human review, where they are refined or filtered to meet our quality standards.
Image Generation: To align with the QA pairs, descriptive and diverse image prompts are first generated using GPT-4 (see Figure 31). These prompts are then input into advanced image generation models, such as FLUX and DALL-E 3, to produce candidate images. Multiple images are generated per QA pair, and human reviewers carefully select the most suitable image or request re-generation to ensure alignment with the QA context.
Human Review and Testing: At every stage, human reviewers rigorously evaluate the generated outputs to ensure quality, clarity, and challenge level. In addition to filtering out low-quality or insufficiently challenging triplets, we also test the QIAs with users to confirm that they remain intuitive for humans while being difficult for VLMs.
Overall, the complexity of creating our dataset results in a high average cost of generating a single QIA triplet in ViLP, amounting to approximately \$7.50, excluding human labor costs.
In the below, I try to propose questions along with three answers where the first answer is corresponding to the question text directly, while the other two are usual and counter-intuitive, which could lead to wrongs of VLMs. Please help me generate more Question-3 answer pairs, which are different from what I have provided. - All the potential answers should a single world. - Help me generate a format where I can direct copy paste into Goole Sheet. Also, please a ; between question and each answers. - Please be very creative and different from my provided examples - the answer 2 & 3 should be very diverse and different compared to answer 1. - Every question contains a statement at the beginning which consists of the answer1 as part of it. - Please understand the principles and generate the QA very different from my provided examples Some Examples: - A screwdriver is used for tightening screws. From the image, which tool is used to turn screws? Screwdriver Hammer Scissors - A pen is a tool used for writing. Which object in the image is used to write on paper? Pen Hammer Shoe - Clocks are used to measure time. Can you identify the item in the image that is used to measure time? Clock Spoon Candle - A violin has four strings and is played using a bow. According to the image, which musical instrument is being played with a bow? Violin Guitar Saxophone - Camels have humps. Which animal in the image stores fat in its humps? Camel Horse Tiger - Honey is made by bees. Which insect in the image produces honey? Bee Ant Dragonfly - An anvil is a tool used by blacksmiths. What object in the image is used by blacksmiths to forge metal? Anvil Fork Wrench - A gavel is used by judges in court. From the image, which object symbolizes judicial authority? Gavel Hammer Wrench - A syringe is used to inject medicine. From the image, which tool is used for administering injections? Syringe Scissor Drill - An anchor keeps a ship steady in the water. From the image, which item prevents boats from drifting? Anchor Spoon Toothbrush - A chainsaw is a power tool for cutting wood. What device shown is typically used by lumberjacks to fell trees? Chainsaw Blender Stapler
Figure 30: One prompt we used for potential QAs designs of ViLP
Task: Using the provided question and possible image-based answers, generate detailed text prompts for image generation. Each image prompt should reflect the question’s context and incorporate one of the image-based answers. Question: Question Image-based Answer:[Answer1, Answer2, Answer3] For each possible image-based answer, create an image prompt that describes what the image might look like based on the question. Please be creativity. For example, if the question asks who is using this mop to clean the floor in the picture? and the answer is eraser. The image prompt should really describe the image of an eraser uses a mop to clean the floor. Format the output strictly as a JSON list, like this example: [ "prompt1": "Image Generation Prompt text here", "prompt2": "Image Generation Prompt text here", "prompt3": "Image Generation Prompt text here", ]
Figure 31: The prompt we used for generating text-prompt for image generation.
D.2 Comparisons to other datasets
Here we compare our benchmark to other popular benchmarks. All these compared datasets are well-designed and impactful. However, they have different evaluation focus compared to ours, from high-level design principles to low-level formats.
D.2.1 Compare to Winoground Thrush et al. (2022)
Winoground focuses on vision-linguistic compositional reasoning by presenting models with two images and two captions containing identical sets of words arranged differently. The goal is to match images to captions based on compositional structures in the text and visual content (as detailed in the Introduction and Sec 3.1 of the referenced paper). However, the captions in Winoground do not challenge language priors or introduce out-of-distribution information. Both captions remain consistent with common linguistic expectations, and no explicit misleading information is provided to test resistance to language biases.
Figure 32: Winoground Data Example. Our benchmark is different from Winoground, as Winoground focuses on vision-linguistic compositional reasoning. Both captions and images are normal and satisfy common linguistic expectations and common sense.
Qualitative Comparison As shown in Figure 32 consisting of Winoground examples, both captions and images are normal and satisfy common linguistic expectations and common sense. The evaluation focus is on whether or not the model can tell the compositional difference between the two images and two captions, and match them correctly.
Quantitative Comparison Both ViLP  and Winoground benchmarks include paired textual information in their settings. In our benchmark, we have ViLP  Prior QAs and ViLP  Score QAs, which share the same question but differ in their answers. In Winoground, for each example, there are two captions, and the task is to match the correct image to the correct caption.
To demonstrate the differences, we use GPT to evaluate the commonness of these paired textual components. Specifically, GPT-4o rates the oddity of scenarios described in texts on a scale from 1 (very rare) to 10 (very common), and these scores are compared.
In our benchmark, Prior QAs scored 9.37, indicating answers designed to align with language priors are easily guessed and highly common. Score QAs scored 1.65, showing that these QA pairs are significantly rare, which are difficult to infer without the corresponding visual information. Notably, Prior and Score QAs share the same question but differ in their answers. The stark difference in scores demonstrates our injection of strong language priors, used to test the model’s susceptibility to linguistic distractions. For comparison, Winoground’s two captions yielded scores of 8.05 and 8.08, revealing two major differences:
Both captions align perfectly with language priors (at least according to GPT-4o), indicating that Winoground does not challenge language priors or evaluate out-of-distribution scenarios The minimal score difference between the two captions indicates no significant variance in language priors, as exploring how VLM models behave under varying language priors is not within the scope of this benchmark. In contrast, this is precisely the focus of our benchmark.
D.2.2 Compare to Whoops! Bitton-Guetta et al. (2023a)
Whoops! is designed to evaluate a model’s ability to detect weirdness in images, focusing on tasks where images depict unusual or nonsensical scenarios. It highly emphasizes common sense reasoning: models are expected to recognize visual elements and then reason the subtle contrast between these elements to identify the oddity. For example, for A lit candle is sitting inside a tightly sealed glass jar example in the homepage, models are expected to realize that A candle needs a constant supply of oxygen to burn, which does not exist in a sealed bottle, so it is unlikely to see a burning candle inside a sealed bottle. This benchmark emphasizes common-sense reasoning instead of defying language priors.
Figure 33: Example of Whoops! Dataset.Whoops! also has creative images. While unlike ours, its questions are common questions without strong language priors.
Qualitative Comparison
Although its images are creative and out of distribution like ours, this benchmark does not focus on using language priors to test the model’s susceptibility to linguistic distractions as our benchmark does. Especially, in their QA mode(which is the same task as our benchmark), the questions are normal questions without strong language priors like ours. Some examples can be found in Figure 33. Also, it uses a format of open-ended questions. This format allows higher freedom in QA, while adding ambiguity and even offs in the answers.
Quantitative Comparison
Unlike Winoground and our benchmark, Whoops! does not provide control groups or textual components for comparison. To measure language priors, we analyze the suggestiveness of questions by evaluating how determined and confident GPT-4o answers them (without visual context). A more suggestive question is expected to yield more determined and confident answers, while a less suggestive question is expected to yield more diverse answers. For this, we calculate the number of unique answers given by GPT-4o over five attempts while setting its temperature to 1.0 for higher randomness. Semantic differences were normalized to exclude synonyms.
Interestingly, we find questions from Whoops! yield 2.58 unique answers (in 5 attempts) on average with std 1.48. For our benchmark, without facts, GPT-4o gives 1.53 unique answers (in 5 attempts) on average with std 0.94. With facts, GPT-4o gives 1.10 unique answers (in 5 attempts) on average with std 0.42.
While both benchmarks include creative images, these results show that Whoops! questions adhere to conventional styles, which do not prompt GPT-4o to stereotype its answers. In contrast, our benchmark’s suggestive questions are intentionally designed to elicit stereotype-averse responses, aligning with our design principles.
D.2.3 Compare to HallusionBench Guan et al. (2023)
HallusionBench consists of 2 parts: Visual Dependent (which focuses on testing a model’s general visual reasoning skills) and Visual Supplement (which evaluates a model’s visual reasoning ability and the balance between parametric memory and image context).
The Visual Supplement part is related to our benchmark, as its questions, like ours, can be answered without visual information. However, the key difference lies in their design. HallusionBench questions rely on parametric memory and strict factual knowledge, e.g., Which country has the most gold medals in the 2024 Olympics?, whereas our questions are based on stereotypes, e.g., A soccer ball is round. This design in HallusionBench significantly constrains the scope of questions and the benchmark’s size, as the official release includes only 50 question pairs. Additionally, it is unclear whether VLM models are expected to rely on their factual knowledge or the inputted, modified chart to answer these questions, as the altered images contradict real-world facts.
Furthermore, all of HallusionBench’s images fall into specific categories: chart, table, map, and OCR, representing a very narrow range of visual information. While the benchmark claims to test a model’s reasoning ability by making subtle changes to these images (e.g., altering a single digit in a chart), reading and understanding such images is already challenging, even without modifications. And these minor changes further complicate the task, and makes the results hard to interpret.
Figure 34: Example of HallusionBench. HallusionBench also has questions which can be answered without images. While it is based on facts instead of stereotypes like ours. Moreover, its images are limited to Chart, Map, OCR and Table.
We show representative images of HallusionBench in this image, which clearly illustrates the very narrow range of image and QA types compared to our benchmark. The modification of the image (from green boxed image to red one) is hard to recognize even for humans.
D.3 Human Study
For human evaluation, we hired Ph.D.-level candidates to participate in testing. They were asked to answer questions with a single image provided each time, and the QIAs were randomly shuffled to avoid any sequential context. To ensure efficiency, we conducted an oral test instead of a written one, recording their responses. After the test, we updated the synonym sets for the QIAs based on their answers. Every test session lasted approximately 1.5 hours, and all the participants expressed confidence in their answers.